# Scalable Multispecies Ion Transport in a Grid Based Surface-Electrode Trap

### v1

@@ III Center-to-left-or-right primitive
The measurements described were all performed on a single pair of ions that serially sampled the numerous wells present in the broadcast transport operations, and all with identical initial and final measurement locations. We validated that the transport operations accomplished their intended outcome in several ways, including a crystal tagging method using Ba-Sr presented in the Sec. D. To clearly demonstrate the ability of the swap-or-stay primitive to sort a fully loaded 2D grid trap in parallel, we loaded Yb ions in all eight legs of the region of interest and implemented a 2D sorting sequence via swap-or-stay operations. A stop motion video showing the trajectory of all Yb ions was collected with a complementary metal-oxide-semiconductor (CMOS) camera (see Supplementary Video S1).
V Reorder-or-not
In addition to swapping qubits between zones, a trapped ion quantum processor may require zone-specific reordering of ion crystals. Such conditional reorders may be part of a larger sorting algorithm [39], or simply be used to correct spontaneous crystal reorders that may occur after a background gas collision [40, 10]. In close analogy with the swap-or-stay sequence, we demonstrate a reorder-or-not operation based on C2LR. This operation is only demonstrated for Yb-Ba, though nothing in principle precludes extending it to Ba-Sr.
This reorder-or-not sequence begins and ends with C2LR, as in swap-or-stay, but with a new intermediate waveform that holds two potential wells but rotates only the left one by 180 degrees (Fig. 5(a)). In Fig. 5(b) we show the resulting excitation of a Yb-Ba crystal following a reorder-or-not operation as a function of angular speed, and as a function of number of round trips in Fig. 5(c). In both Fig. 5(b-c) we average the resulting occupancy over all six horizontal zones of the trap as well as both conditions. The error bars represent one standard deviation on the average value of the excitation across all measured zones.
From the multiple reorder-or-not operations performed sequentially, we can infer heating rates per operation and compare this to baseline rates. For C2LR transport speed of 10 m/s and an angular speed of rad/ms, we infer quanta/reorder-or-not in the XCOM mode, and quanta/reorder-or-not for the XSTR mode. Including interpolation time, reorder-or-not takes 93 µs to complete, and with our measured XCOM static heating rate of 181 quanta/s predicts quanta/reorder-or-not (see Tab. 3 for heating rate data). The XSTR mode heats less at 12.3 quanta/s, giving quanta/reorder-or-not. As in the case of swap-or-stay, the heating inferred during transport is close to that of our background heating rates seen in static potential wells with approximately equivalent axial frequencies.
@@ VI Conclusion
The success of the transport operations, attaining sub-quanta motional excitation at 3 m/s speeds across multiple sites in the trap and with crystals of different composition, confirms the necessary control for multispecies junction transport can be realized despite the restrictions imposed by extensive electrode co-wiring. Looking forward, scaling to a larger grid should improve transport performance and simplify waveform development as the trap will more closely resemble a periodic tiling.
While here the voltages for the different C2LR conditions were exchanged in software, that control can readily be moved in vacuum [41] and even as far downstream as the trap chip itself through the delivery of binary signals and integrated double-pole single-throw switches. Compared to analog signals, the delivery of a large number of binary signals presents minimal technical challenge, and draws directly from application-specific integrated circuit techniques used in classical computing. The number of on-chip switches needed to implement this method for sorting qubits in a large scale trap are well within the capabilities of cryogenic CMOS technology, and indeed this is already an active area of research for ion traps [42, 43].
@@ Acknowledgements.
- Clark et al. [2021] C. R. Clark, H. N. Tinkey, B. C. Sawyer, A. M. Meier, K. A. Burkhardt, C. M. Seck, C. M. Shappert, N. D. Guise, C. E. Volin, S. D. Fallek, et al., High-fidelity bell-state preparation with Ca optical qubits, Physical Review Letters 127, 130505 (2021).
- Harty et al. [2014] T. P. Harty, D. T. Allcock, C. J. Ballance, L. Guidoni, H. A. Janacek, N. M. Linke, D. N. Stacey, and D. M. Lucas, High-fidelity preparation, gates, memory, and readout of a trapped-ion quantum bit, Physical review letters 113, 220501 (2014).
@@
- Malinowski et al. [2023] M. Malinowski, D. Allcock, and C. Ballance, How to wire a -qubit trapped-ion quantum computer, PRX Quantum 4, 040313 (2023).
@@
- Blakestad et al. [2011] R. Blakestad, C. Ospelkaus, A. VanDevender, J. Wesenberg, M. Biercuk, D. Leibfried, and D. J. Wineland, Near-ground-state transport of trapped-ion qubits through a multidimensional array, Physical Review A 84, 032314 (2011).
@@
- Shu et al. [2014] G. Shu, G. Vittorini, A. Buikema, C. Nichols, C. Volin, D. Stick, and K. R. Brown, Heating rates and ion-motion control in a Y-junction surface-electrode trap, Physical Review A 89, 062308 (2014).
@@
- Barrett et al. [2003] M. D. Barrett, B. DeMarco, T. Schaetz, V. Meyer, D. Leibfried, J. Britton, J. Chiaverini, W. Itano, B. Jelenković, J. Jost, et al., Sympathetic cooling of and for quantum logic, Physical Review A 68, 042302 (2003).
@@
- Lancellotti et al. [2023] F. Lancellotti, S. Welte, M. Simoni, C. Mordini, T. Behrle, B. de Neeve, M. Marinelli, V. Negnevitsky, and J. Home, Low-excitation transport and separation of high-mass-ratio mixed-species ion chains, arXiv preprint arXiv:2312.14006 (2023).
- Ruster et al. [2014] T. Ruster, C. Warschburger, H. Kaufmann, C. T. Schmiegelow, A. Walther, M. Hettrich, A. Pfister, V. Kaushal, F. Schmidt-Kaler, and U. G. Poschinger, Experimental realization of fast ion separation in segmented Paul traps, Physical Review A 90, 033410 (2014).
- Murali et al. [2020] P. Murali, D. M. Debroy, K. R. Brown, and M. Martonosi, Architecting noisy intermediate-scale trapped ion quantum computers, in 2020 ACM/IEEE 47th Annual International Symposium on Computer Architecture (ISCA) (IEEE, 2020) pp. 529–542.
@@
- Rasmusson et al. [2021] A. Rasmusson, M. D’Onofrio, Y. Xie, J. Cui, and P. Richerme, Optimized pulsed sideband cooling and enhanced thermometry of trapped ions, Physical Review A 104, 043108 (2021).
@@
- Hankin et al. [2019] A. M. Hankin, E. Clements, Y. Huang, S. M. Brewer, J.-S. Chen, C.-W. Chou, D. Hume, and D. Leibrandt, Systematic uncertainty due to background-gas collisions in trapped-ion optical clocks, Physical Review A 100, 033419 (2019).
@@
- Stuart et al. [2019] J. Stuart, R. Panock, C. Bruzewicz, J. Sedlacek, R. McConnell, I. Chuang, J. Sage, and J. Chiaverini, Chip-integrated voltage sources for control of trapped ions, Physical Review Applied 11, 024010 (2019).
@@ Appendix A Effective heating rate during transport
For slow transport speeds ( kHz), we observe a linear increase in excitation with increased transport time that can be described by an effective heating rate. The linear fits are plotted in Fig. 6 and the resulting effective heating rates are shown in Tab. 3 alongside the measured heating rates in the zone centers for Yb-Ba (Ba-Sr) in multiple horizontal (vertical) zones.
For both Ba-Sr and Yb-Ba, the effective heating rates of the XCOM modes are comparable to the heating rates measured in a static well at the zone center (Tab. 3), while XSTR heats significantly more during a swap than a stay. This good XCOM agreement is despite the complicated behavior that occurs as ions approach the junction, involving excursions of the crystal off of the pseudopotential minima, rapidly changing motional mode frequencies and directions, including population exchange between different normal modes.
@@ Appendix C Static potential well frequencies
A key component of our broadcast electrode scheme is the uniformity/periodicity of the trap zones. One experimentally accessible proxy for this uniformity, in addition to the success of transport operations in different legs presented in the main text, is the consistency of the motional mode frequencies from leg to leg (using Yb-Ba). These frequencies are presented in six legs for Yb-Ba in Tab. 4.
Appendix D Validation of swap-or-stay transport operations
As we can run transport with the Yb-Ba (Ba-Sr) crystals remaining in their quantum ground state, and that this agrees with our numerical simulations, is sufficient evidence that the swap/stay is being completed as expected.
To directly verify the intended outcome of the operation in parallel operation, two Ba-Sr crystals were loaded into adjacent horizontal and vertical zones. The two crystals are Doppler cooled using a pair of co-propagating 422 nm and 493 nm beams aligned at 45 degrees to the grid axis, illuminating both crystals simultaneously, with repump beams (at 1092 nm and 650 nm) delivered as sheet beams. A single 1762 nm beam aligned with the vertical leg drives transitions between the to states of the Ba ion. The state is isolated from the fluorescence cycling transitions, and so the ion is dark if driven to this ’shelved’ state [5].
@@ Appendix D Validation of swap-or-stay transport operations
A third experimental setup and nominally identical trap went farther and demonstrated basic sorting using the swap-or-stay operation. A 369 nm sheet beam was used to address single ions in a fully loaded section of a section of the grid trap, and the resulting fluorescence was detected by a CMOS camera spanning the entire grid section of the trap. In Supplementary Video S1, using the exact same voltage waveforms for the Yb-Ba data shown in the main text, we show swapping of Yb ions between the vertical zones and horizontal zones above and to the left. We chose to record the video using a third experimental setup as this system had a sheet beam and camera setup that was capable of imaging the entire grid section of the trap. Single Yb atoms were used since a Ba sheet beam was not immediately available.

### v2

@@ III Center-to-left-or-right primitive
All data presented in the main text are taken using a single multispecies ion crystal (Yb-Ba or Ba-Sr) that serially samples the numerous wells present in the transport operations and all with identical initial and final measurement locations. As the Coulomb interaction between ions in different wells is small (Appendix D.4), such serial testing is an effective proxy for performance of a sort in a densely occupied configuration. Additionally, we validate that the transport operations accomplish their intended outcome in several ways, including a crystal tagging method using Ba-Sr (Appendix D). And finally, we visually demonstrate the ability of the swap or stay primitive to sort a fully loaded 2D grid trap in parallel. We load Yb ions in all eight legs of the region of interest and implement a 2D sorting sequence via swap or stay operations. A stop motion video showing the trajectory of all Yb ions was collected with a complementary metal-oxide-semiconductor (CMOS) camera (see Supplementary Video S1).
V Reorder or not
In addition to swapping qubits between zones, a trapped ion quantum processor may require zone-specific reordering of ion crystals. Such conditional reorders may be part of a larger sorting algorithm [37], or simply be used to correct spontaneous crystal reorders that may occur after a background gas collision [38, 10]. In close analogy with the swap or stay sequence, we demonstrate a reorder or not operation based on C2LR. This operation is demonstrated only for Yb-Ba, though nothing, in principle, precludes extending it to Ba-Sr.
This reorder or not sequence begins and ends with C2LR, as in swap or stay, but with a new intermediate waveform that holds two potential wells but rotates only the left one by (Fig. 5(a)). In Fig. 5(b) we show the resulting excitation of a Yb-Ba crystal following a reorder or not operation as a function of angular speed, and as a function of number of round-trips in Fig. 5(c). In both Fig. 5(b-c) we average the resulting occupancy over all six horizontal zones of the trap as well as both conditions. The error bars represent one standard deviation on the average value of the excitation across all measured zones.
From the multiple reorder or not operations performed sequentially, we can infer heating rates per operation and compare this to baseline rates. For C2LR transport speed of 10 m/s and an angular speed of rad/ms, we infer quanta/reorder or not in the XCOM mode, and quanta/reorder or not for the XSTR mode. Including interpolation time, reorder or not takes 93 µs to complete, and with our measured XCOM static heating rate of 181 quanta/s predicts approximately quanta/reorder or not (see Table 3 for heating rate data). The XSTR mode heats less at 12.3 quanta/s, giving approximately quanta/reorder or not. As in the case of swap or stay, the heating inferred during transport is close to that of our background heating rates seen in static potential wells with approximately equivalent axial frequencies.
@@ VI Conclusion
The success of the transport operations, attaining subquanta motional excitation at 3 m/s speeds across multiple sites in the trap and with crystals of different composition, confirms the necessary control for multispecies junction transport can be realized despite the restrictions imposed by extensive electrode cowiring. Looking forward, scaling to a larger grid should improve transport performance and simplify waveform development, as the trap will more closely resemble a periodic tiling.
While here the voltages for the different C2LR conditions were exchanged in software, that control can readily be moved in vacuum [39] and even as far downstream as the trap chip itself through the delivery of binary signals and integrated double-pole single-throw switches. Compared to analog signals, the delivery of a large number of binary signals presents minimal technical challenge, and draws directly from application-specific integrated circuit techniques used in classical computing. The number of on-chip switches needed to implement this method for sorting qubits in a large-scale trap are well within the capabilities of cryogenic CMOS technology, and indeed this is already an active area of research for ion traps [40, 41].
@@ Acknowledgements.
- Clark et al. [2021] C. R. Clark, H. N. Tinkey, B. C. Sawyer, A. M. Meier, K. A. Burkhardt, C. M. Seck, C. M. Shappert, N. D. Guise, C. E. Volin, S. D. Fallek, et al., High-fidelity bell-state preparation with 40Ca+ optical qubits, Physical Review Letters 127, 130505 (2021).
- Harty et al. [2014] T. P. Harty, D. T. C. Allcock, C. J. Ballance, L. Guidoni, H. A. Janacek, N. M. Linke, D. N. Stacey, and D. M. Lucas, High-fidelity preparation, gates, memory, and readout of a trapped-ion quantum bit, Physical Review Letters 113, 220501 (2014).
@@
- Malinowski et al. [2023] M. Malinowski, D. T. C. Allcock, and C. J. Ballance, How to wire a -qubit trapped-ion quantum computer, PRX Quantum 4, 040313 (2023).
@@
- Blakestad et al. [2011] R. B. Blakestad, C. Ospelkaus, A. P. VanDevender, J. H. Wesenberg, M. J. Biercuk, D. Leibfried, and D. J. Wineland, Near-ground-state transport of trapped-ion qubits through a multidimensional array, Physical Review A 84, 032314 (2011).
@@
- Shu et al. [2014] G. Shu, G. Vittorini, A. Buikema, C. S. Nichols, C. Volin, D. Stick, and K. R. Brown, Heating rates and ion-motion control in a Y-junction surface-electrode trap, Physical Review A 89, 062308 (2014).
@@
- Barrett et al. [2003] M. D. Barrett, B. DeMarco, T. Schaetz, V. Meyer, D. Leibfried, J. Britton, J. Chiaverini, W. M. Itano, B. Jelenković, J. D. Jost, C. Langer, T. Rosenband, and D. J. Wineland, Sympathetic cooling of and for quantum logic, Physical Review A 68, 042302 (2003).
@@
- Deen et al. [2024] D. Deen, G. Vittorini, and N. Burdick, Ion trap apparatus with integrated switching apparatus (2024), US Patent 11,876,092.
@@
- Rasmusson et al. [2021] A. J. Rasmusson, M. D’Onofrio, Y. Xie, J. Cui, and P. Richerme, Optimized pulsed sideband cooling and enhanced thermometry of trapped ions, Physical Review A 104, 043108 (2021).
@@
- Hankin et al. [2019] A. M. Hankin, E. R. Clements, Y. Huang, S. M. Brewer, J.-S. Chen, C.-W. Chou, D. B. Hume, and D. R. Leibrandt, Systematic uncertainty due to background-gas collisions in trapped-ion optical clocks, Physical Review A 100, 033419 (2019).
@@
- Stuart et al. [2019] J. Stuart, R. Panock, C. D. Bruzewicz, J. A. Sedlacek, R. McConnell, I. L. Chuang, J. M. Sage, and J. Chiaverini, Chip-integrated voltage sources for control of trapped ions, Physical Review Applied 11, 024010 (2019).
@@
- Doret et al. [2012] S. C. Doret, J. M. Amini, K. Wright, C. Volin, T. Killian, A. Ozakin, D. Denison, H. Hayden, C. Pai, R. E. Slusher, et al., Controlling trapping potentials and stray electric fields in a microfabricated ion trap through design and compensation, New Journal of Physics 14, 073012 (2012).
@@ Appendix A Effective heating rate during transport
For slow transport speeds ( kHz), we observe a linear increase in excitation with increased transport time that can be described by an effective heating rate. The linear fits are plotted in Fig. 6 and the resulting effective heating rates are shown in Table 3 alongside the measured heating rates in the zone centers for Yb-Ba (Ba-Sr) in multiple horizontal (vertical) zones.
For both Ba-Sr and Yb-Ba, the effective heating rates of the XCOM modes are comparable to the heating rates measured in a static well at the zone center (Table 3), while XSTR heats significantly more during a swap than a stay. This good XCOM agreement is despite the complicated behavior that occurs as ions approach the junction, involving excursions of the crystal off of the pseudopotential minima, rapidly changing motional mode frequencies and directions, including population exchange between different normal modes.
@@ Appendix C Static potential well frequencies
A key component of our broadcast electrode scheme is the uniformity/periodicity of the trap zones. One experimentally accessible proxy for this uniformity, in addition to the success of transport operations in different legs presented in the main text, is the consistency of the motional mode frequencies from leg to leg (using Yb-Ba). These frequencies are presented in six legs for Yb-Ba in Table 4.
Appendix D Validation of swap or stay transport operations
D.1 Overview
Ultimately, grid-based QCCD devices will need to be run with all horizontal and vertical legs filled with ions, but a discussion of the required sorting algorithms and hardware for such a device is beyond the scope of this work. All transport operations presented in the main text are demonstrated using a single Yb-Ba or Ba-Sr crystal, and here we discuss various validation experiments to confirm that the ions are undergoing the correct swap or stay transport operations.
It first should be pointed out that we can run transport with the Yb-Ba (Ba-Sr) crystals remaining in approximately their quantum ground state following a swap or stay operation, and that this agrees well with our numerical simulations. This already provides strong evidence that the swap or stay is being completed as expected, but we demonstrate additional validation steps below.
D.2 Parallel Ba-Sr ion crystal swapping
To directly verify the intended outcome of the operation in parallel operation, two Ba-Sr crystals were load into adjacent horizontal and vertical zones. The two crystals are Doppler cooled using a pair of copropagating 422 nm and 493 nm beams aligned at 45 degrees to the grid axis, illuminating both crystals simultaneously, with repump beams (at 1092 nm and 650 nm) delivered as sheet beams. A single 1762 nm beam aligned with the vertical leg drives transitions between the to states of the Ba ion. The state is isolated from the fluorescence cycling transitions, and so the ion is dark if driven to this ’shelved’ state [5].
@@ D.2 Parallel Ba-Sr ion crystal swapping
D.3 Single Yb transport interrogated by a Doppler sheet beam
A third experimental setup and nominally identical trap goes farther and demonstrates basic sorting using the swap or stay operation. A 369 nm sheet beam is used to address single ions in a fully loaded section of a section of the grid trap, and the resulting fluorescence was detected by a CMOS camera spanning the entire grid section of the trap. In Supplementary Video S1, using the exact same voltage waveforms for the Yb-Ba data shown in the main text, we show swapping of Yb ions between the vertical zones and horizontal zones above and to the left. We choose to record the video using a third experimental setup as this system had a sheet beam and camera setup that was capable of imaging the entire grid section of the trap. Single Yb ions are used since a Ba sheet beam was not immediately available.
D.4 Stray fields from other ion crystals
As shown in Fig. 3, during a swap or stay operation the ions are continually shuttled into and out of the junction during the swap procedure. This leaves open the possibility of Coulomb interactions between neighbouring ion crystals affecting transport. To bound the effects of this interaction we can use the minimum separation between two ion crystals, which occurs when one crystal is in the junction and the other is in a static horizontal or vertical well–see for example Fig. 3, step (e). In this configuration the separation between the two crystals is and stray electric field then consists of the direct Coulomb interaction, and the Coulomb field resulting from the effective image charge generated by the trap top metal
For and the magnitude of the stray field is V/m, where is the electron charge. As stray fields from surface imperfections, local charging and other effects are typically in the range of 10-100 V/m in surface-electrode traps [49], the effective stray field due to Coulomb potentials can safely be ignored, even in a densely loaded grid. This indicates that measurements of transport of a single multispecies ion crystal throughout the trap faithfully represents the motional heating one would expect to measure in a fully loaded trap.